### v1

# Bin Yang1,2, Alexandru Paul Condurache1,2

Abstract
3D scene understanding is a critical yet challenging task in autonomous driving, primarily due to the irregularity and sparsity of LiDAR data, as well as the computational demands of processing large-scale point clouds. Recent methods leverage the range-view representation to improve processing efficiency. To mitigate the performance drop caused by information loss inherent to the ”many-to-one” problem, where multiple nearby 3D points are mapped to the same 2D grids and only the closest is retained, prior works tend to choose a higher azimuth resolution for range-view projection. However, this can bring the drawback of reducing the proportion of pixels that carry information and heavier computation within the network. We argue that it is not the optimal solution and show that, in contrast, decreasing the resolution is more advantageous in both efficiency and accuracy. In this work, we present a comprehensive re-design of the workflow for range-view-based LiDAR semantic segmentation. Our approach addresses data representation, augmentation, and post-processing methods for improvements. Through extensive experiments on two public datasets, we demonstrate that our pipeline significantly enhances the performance of various network architectures over their baselines, paving the way for more effective LiDAR-based perception in autonomous systems. The code will be released based on the acceptance.
1 Introduction
LiDAR is one of the most common sensors for perception in autonomous driving. Semantic segmentation on LiDAR point clouds is essential for getting useful and reliable information of the surrounding 3D environment. To solve this 3D scene understanding task, many prior works propose to integrate deep learning techniques because of its remarkable advancements in the past few years. The publication of various annotated datasets [11, 3, 30] in the domain of autonomous driving further promotes research in the field.
In general, those methods can be categorized based on LiDAR data representation into point-based [47, 26, 39], voxel-based [49, 14] and projection-based methods [46, 24, 5, 48]. Both point- and voxel-based approaches typically require substantial computational resources due to the need to process data through networks with numerous 3D convolutional layers, intensive feature pre-processing, and deep architectures involving multiple downsampling and upsampling operations. These requirements can result in slow inference speeds, limiting their suitability for real-time applications. In contrast, rasterizing point cloud into range-view images [7] is more advantageous in fast and scalable LiDAR perception, because it allows the use of 2D operators for efficient computation and facilitates the transfer of knowledge from camera images [1, 18].
[…]
Nevertheless, learning from range-view representations can suffer from the performance drop caused by the ”many-to-one” conflict of adjacent points. To resolve the problem, most approaches maximize the azimuth resolution to make the range image more informative [24, 7, 5, 48, 13, 18]. We argue that this is suboptimal for the following reasons: 1) High overhead: increasing resolution significantly raises computational demand, limiting real-time performance. 2) Inefficient computation: the sparse nature of LiDAR data leads to many empty grid cells. Those unoccupied pixels introduce noise in the data and consume resources needlessly. 3) Limited informative gains: due to distortions caused by ego-motion and sensor-internal errors, points mapped to the same azimuthal location do not perfectly align with the laser beams. This misalignment implies that prioritizing both azimuth and elevation resolutions, rather than increasing image width alone, would yield a higher projection rate, as seen in our analysis in Fig. 2.
Figure 2: Statistics on SemanticKITTI [2]: 3D validity (proportion of projected points) with different azimuth (W) and elevation (H) resolutions. Comparable increases are observable when doubling azimuth and elevation resolution ($\Delta V_{azi},\Delta V_{ele}$).
Building on the aforementioned insights, we introduce FLARES, a novel training schema for range-view semantic segmentation of LiDAR point clouds. Rather than focusing on designing a new network architecture, FLARES targets to address the common limitations outlined above, hence, our proposal is generalizable to any range-view-based approaches. In essence, FLARES divides the point cloud into multiple sub-clouds, with only one projected onto a low-resolution image during training. During inference, all sub-clouds are projected and stacked along the batch dimension for processing. This approach increases the projection rate in both azimuth and elevation, thereby enriching informativeness of the range-view representation at lower cost.
Downscaling resolution, however, can exacerbate class imbalance in the dataset [7], potentially leading to overfitting during training. Another problem is decreasing 2D occupancy due to splitting the point cloud. To tackle these two incidental issues, we extend the pipeline with two additional data augmentation steps. Furthermore, we explore improvements in post-processing methods, a topic that has received minimal attention in previous works.
In summary, our contributions are as follows: 1) We introduce FLARES, a newly designed training schema for faster and more accurate LiDAR semantic segmentation. 2) We integrate two data augmentation techniques tailored to the new schema to enhance the network performance. 3) We propose a novel interpolation-based post-processing approach to resolve the ”many-to-one” problem more effectively. 4) We generalize this schema and all components across various range-view semantic segmentation networks, demonstrating superior performance over baselines on two different benchmarks.
2 Related Works
Point- and Voxel-based methods Some recent works [26, 47, 38] use raw point cloud data as direct network input, eliminating the need for post-processing after prediction. However, these methods often face high computational complexity and memory usage. To address these issues, Hu et al. [17] introduced sub-sampling and feature aggregation techniques for large-scale point clouds to reduce computational costs. Despite these efforts, performance degradation remains significant. Other works [49, 19] use 3D voxel grids as input, achieving point-based accuracy with reduced computational costs by utilizing sparse 3D convolutions [6]. Nonetheless, voxelization and de-voxelization steps continue to be time- and memory-intensive. Range-view-based methods To address inefficiencies, some prior works [36, 24, 37] convert the large-scale point cloud to panoramic range image through spherical projection and leverage image segmentation techniques for LiDAR data. SalsaNext [7] uses a Unet-like network with dilated convolutions to broaden receptive fields for more accurate segmentation, while Lite-HDSeg [28] introduces an efficient framework using a lite version of harmonic convolutions. Additionally, FIDNet [48] and CENet  [5] interpolate and concatenate multi-scale features with a minimal decoder for semantic prediction. These methods share the benefit of lightweight network design, significantly improving efficiency and enabling real-time applications. Nevertheless, they generally underperform 3D methods due to the “many-to-one” issue, where multiple points project to the same pixel. To offset the performance drop caused by the problem, some other recent works propose to use Vision Transformer (ViT) [9, 41, 35]. RangeViT [1] deploys standard ViT backbone as encoder, followed by a light-weight decoder for refining the coarse patch-wise ViT representations, while RangeFormer [18] utilizes a pyramid-wise ViT-encoder to extract multi-scale features from range images. ViTs offer higher model capacities and excel at capturing long-range dependencies by modeling global interactions between different regions, enhancing segmentation performance over traditional CNNs [8]. However, the quadratic computational complexity of self-attention mechanisms in ViTs introduces challenges in achieving an optimal balance between efficiency and accuracy. Training schema Highlighting inefficiencies stemming from the use of high-resolution range images, some methods have adopted compact networks, significantly reducing network capacity [48, 7, 5]. Unfortunately, high-resolution range images still demand substantial memory, which restricts scalability in terms of batch size and data throughput. Lowering the resolution exacerbates information loss, leading to inferior results. To address this, Kong et al. [18] proposed Scalable Training from Range view (STR), a strategy that divides range images into multiple sub-images from different perspectives to reduce memory consumption. Although STR lessens memory usage, it results in a slight drop in segmentation accuracy and only minor improvement in inference speed compared to the baseline. Augmentation Data augmentation plays a crucial role in helping models learn more generalized representations, thereby enhancing scalability. For example, Mix3D [25] introduced an out-of-context mixing strategy by fusing two scenes. Similarly, MaskRange [13] proposed a weighted paste-drop augmentation to manually balance the class frequencies, while RangeFormer [18] employed four consecutive range-wise operations to provide richer semantic and structural cues in the scene. In this work, we introduce two novel augmentation steps, including point-wise and range-wise fusion, specifically designed for the range images utilized in our new schema. Post-Processing Addressing the prevalent ”many-to-one” problem in range-view representations often necessitates a post-processing step to upsample 2D predictions, a critical yet underexplored area in prior research. For example, RangeViT [1] introduced a trainable 3D refiner using KPConv [32]. However, while it directly optimizes 3D semantics, the performance improvement is limited, and the approach adds significant computational overhead. Some methods rely on conventional unsupervised techniques to infer semantics for 3D points; for instance, Milioto et al. [24] proposed a KNN-based voting approach, and Zhao et al. [48] introduced Nearest Label Assignment (NLA), which assigns labels based on the closest labeled point in 3D space. Nevertheless, these unsupervised techniques often struggle with accurately predicting boundaries and distant points, with performance further declining as range-image resolution decreases. To overcome these limitations, we design a new post-processing method that better interpolates the predictions of unprojected points using neighborhood information.
3 Proposed Approach
Our method introduces critical optimizations across three areas: a distinct training and inference scheme for low-resolution range images, advanced augmentation techniques, and an effective post-processing approach for accurately mapping 2D predictions into 3D space.
3.1 Pre-processing
Range-view Projection A LiDAR point cloud consists of points captured during a single revolution, denoted as $P=\{p_{1},...,p_{n}\}$, where each measurement represents a 4D point including the cartesian coordinates $p_{i}=\{x_{i},y_{i},z_{i}\}$ and intensity $t_{i}$. By pre-defining the 2D resolution, assuming $W$ and $H$ as the image width and height, we can project the point cloud into a range image, where each row $v$ and column $u$ correspond to the elevation and azimuth angles of LiDAR points. The mathematical expression of the spherical projection model is as following:
$\begin{bmatrix}u\\ v\end{bmatrix}=\begin{bmatrix}\frac{W}{2}-\frac{W}{2\pi}arctan(\frac{y}{x})\\ \frac{H}{\Theta_{max}-\Theta_{min}}*(\Theta_{max}-arcsin(\frac{z}{d}))\end{bmatrix}$ (1)
Angular values $\Theta_{max}$ and $\Theta_{min}$ define the upper and lower bound of the LiDAR’s vertical field of views and the depth value is calculated by $d=\sqrt{x^{2}+y^{2}+z^{2}}$. Note that $H$ is typically determined by the number of LiDAR sensor beams, while W can be assigned with random value based on the requirements. Similar to the prior studies [5, 48, 7], we adopt a five-channel input representation ($x,y,z,t,d$). Data Representation Increasing azimuth resolution can help preserve more details from point clouds, but this comes at the expense of reduced efficiency. Additionally, as illustrated in Fig. 2, elevation resolution plays an equally important role in maximizing projection rates. This led us to rethink if a large image width is necessary to mitigate information loss. Inspired by STR [18], we propose an alternative solution: downsampling the point cloud into multiple equal-interval sub-clouds and projecting them into range images with a lower azimuth resolution. During training, we randomly select one range image. For inference, we stack all images along the batch dimension and process them in a single forward pass. We depict the workflow of FLARES in Fig. 3. To balance 3D validity and 2D occupancy, assuming $N_{max}$ is the maximum number of partitioning groups for the specific resolution of range images, it is determined by the rule that the average 2D occupancy must not fall below the high-resolution range image in standard mode ($\frac{1}{N_{max}}\sum^{N_{max}}_{i}Occ_{i}\geq Occ_{high}$). This new design offers three key advantages: 1) Enhanced projection rate by increasing both image height and width; 2) Reduced memory consumption, enabling deployment on smaller GPUs; 3) Preservation of the full field of view, maintaining contextual integrity despite downsampling.
Figure 3: Visual Illustration of FLARES: The full LiDAR point cloud is equally divided and grouped, each projected into a lower-resolution range image. During training, one image and its 2D label are sampled for optimization. For testing, stacked sub-cloud projections are processed simultaneously, and the outputs are fused into 3D predictions using an unsupervised method. We provide more details in the supplementary material.
3.2 Data Augmentation
Previous studies [5, 1, 48] have primarily used geometric transformations such as random flipping, translation, and rotation for point cloud augmentation. To further enrich semantic contexts and optimize for the low-resolution range image used in FLARES, we introduce two additional augmentation steps: 1) Weighted Paste-Drop+ Class imbalance is a common issue in LiDAR semantic segmentation benchmarks [2, 3], where certain classes are heavily underrepresented. This imbalance is compounded by information loss during point cloud-to-range image conversion, especially when downscaling azimuth resolution. Building upon the Weighted Paste Drop (WPD) approach introduced in MaskRange [13], which pastes pixels from rare classes and drops pixels from abundant classes, we present an enhanced version, WPD+. Unlike the original method, which performs geometric data augmentation identically on both sampled and current frames in 3D space before projection, our approach applies WPD directly in 3D space, which avoids the repeating computation of geometric transformations, and samples multiple frames to improve class balancing. Additionally, we use a small set of synthetic dataset generated in the Carla Simulator [10] to further augment rare classes that correspond to small and dynamic objects in the scene. Despite possible domain gaps between datasets, it yields notable accuracy improvements from our experimental results. 2) Multi-Cloud Fusion Given the inherent sparsity of LiDAR point clouds, range images often contain a considerable number of empty grids. Prior works have mitigated the issue by interpolating missing pixels from nearest neighbors [44] or by supplementing occupied pixels from other frames [18]. Coming down to the point cloud splitting in FLARES mode, the 2D occupancy in the single range image is decreased and it prompts the necessity for some solution. In our empirical study in Fig. 5(d), we found that it yields sub-optimal results when training directly on sub-cloud. To address the issue, we propose Multi-Cloud Fusion (MCF), a strategy to increase the 2D occupancy of sub-clouds alongside the setup of FLARES. Assuming the point cloud can be divided into a maximum of $N_{max}$ groups, we randomly pick a group number $N$in $\{1,2,...,N_{max}\}$ and split the full point cloud accordingly. After projecting them into $N$ range images, we randomly select one $R\in\mathbb{R}^{H\times W\times 5}$ as the training input. To enhance occupancy, all empty pixels in $R$ are filled using occupied pixels from remaining $N-1$ range images. This method maximizes the 2D occupancy in the range image of a sub-cloud while maintaining the structural consistency of the scene.
For further technical details of how input data is curated and augmented, please refer to the supplementary material.
3.3 Post-Processing
After the augmented image being processed, 2D predictions from the network must be reprojected into 3D space using some post-processing technique. To align with the new inference framework with stacked predictions, we first propose an extension of the standard KNN method [24], termed  KNN Ensembling. In the post-processing phase, all sub-clouds are iteratively processed with KNN and votes are ensembled for every point from the full cloud to obtain final predictions. However, we found that the extension still faces the challenge in appropriately weighting contributions of nearest neighbors in 3D coordinates. In addition, the inference time is accumulated due to iterative process. To address the limitation, we propose a novel algorithm called Nearest Neighbors Range Interpolation (NNRI). Its pseudo-code is detailed in Algo. 1.
After applying softmax to the network output, we begin by kernelizing 2D predictions and range images using a pre-defined kernel size ($3\times 3$ in our experiments). Next, we assign each point’s nearest neighbors in 2D space with corresponding 2D coordinates and stack them along the sub-cloud dimension. The relative depth between each point and its neighbors is computed by taking the absolute difference in depth values. To extract valid data for interpolation, a threshold is needed to filter out distant neighbors. According to the prior knowledge [16, 20], using a constant threshold is sub-optimal due to differing point densities in LiDAR data: closer points are more likely to be affected by outliers due to high density, while farther points struggle to find valid neighbors due to sparsity. To fit this underlying geometry, the range value of each point is employed to determine its cut-off value. By normalizing the range using pre-computed mean and standard deviation, the cut-off value is derived from an exponential function, which approximates the relationship between point-sensor distance and density [21]. This approach simplifies computation by avoiding the costly nearest neighbor search in 3D space to calculate exact density values and adaptively assigns a threshold to each point. Once valid nearest neighbors are identified, they are normalized within the range of $[0,1]$ to compute interpolation weights. Finally, softmax scores of all 3D points are interpolated by the weighted sum of their nearest neighbors. NNRI is designed to effectively mitigate the ”many-to-one” issue inherently in range-view methods by leveraging distance-wise local neighborhood information in both 2D and 3D.
Define : $N=N_{max}$ sub-clouds.
The annotation contains $C$ classes
Input : Range images $R_{ranges}$ with size $N\times H\times W$,
Softmax scores $I_{scores}$ with size $N\times C\times H\times W$,
Arrays $R_{all}(p)$ with range values for all points,
Image coordinates $(u_{all},v_{all})$ for all points,
Kernel size $k$,
Padding $pad$,
Cut-off factor $\alpha$,
Mean of all range values $r_{mean}$,
Standard Deviation of all range values $r_{std}$
Output : Array $Labels$ with predicted labels for all points.
1:Unfold scores and ranges with $k\times k$ kernel:
2:    $S_{s}(n,h,w,k)\leftarrow\texttt{unfold}(I_{scores},k,pad)$
3:    $S_{r}(n,h,w,k)\leftarrow\texttt{unfold}(R_{ranges},k,pad)$
4:Extract nearest-neighbors for each point $p$:
5:    $N_{s}(n,p,k)\leftarrow S_{s}(n,h,w,k)[...,u_{all},v_{all}]$
6:    $N_{r}(n,p,k)\leftarrow S_{r}(n,h,w,k)[...,u_{all},v_{all}]$
7:Compute relative depths:
8:    $N_{rel}(n,p,k)\leftarrow||(N_{r}(n,p,k)-R_{all}(p)||$
9:Compute the cut-off value for each point $p$:
10:    $D(p)=\texttt{exp}(\frac{R(p)-r_{mean}}{r_{std}})*\alpha$
11:Filter the valid neighbors and compute weights:
12: $N_{valid}(n,p,k)\leftarrow\texttt{clamp}(N_{rel}(n,p,k),\texttt{max}=D(p)$)
13:    $W(n,p,k)=1-\texttt{Normalize}(N_{valid}(n,p,k))$
14:Weighted Sum for 3D Projection:
15:    $Scores(p)=\sum_{i}^{k^{2}\times n}W(n,p,k)*N_{s}(n,p,k)$
16:    $Labels=\texttt{argmax}_{c\in C}(Scores(p))$
17:Return $Labels$
Algorithm 1 Nearest Neighbors Range Interpolation
3.4 Network Selection
In order to pursue enhancement in the segmentation accuracy while possibly maintaining the high efficiency of range-view-based approaches, we revisited prior works and selected three light-weight CNN-based networks (FIDNet [48], SalsaNext [7] and CENet [5]) for integration. To further test the effectiveness across different network architectures, we additionally deploy RangeViT [1], a network composed of a series of Vision Transformer blocks [9], in the experimental phase. Original RangeViT uses a trainable KPConv-based 3D projector to get the point-wise predictions. We replace it with our post-processing component to achieve the full integration of our framework and train the model from scratch.
4 Experimental Analysis
4.1 Settings
Datasets We conduct experiments on two public LiDAR semantic segmentation datasets. SemanticKITTI [2] dataset [2] consists of 22 sequences captured with a 64-beam LiDAR sensor, encompassing 19 semantic classes. The dataset is split as follows: sequences 00 to 10 (excluding 08) are used for training, sequence 08 is reserved for validation, and sequences 11 to 21 are designated for testing. nuScenes dataset [3] comprises 1,000 driving scenes recorded in Boston and Singapore using a 32-beam LiDAR sensor, leading to a relatively sparse point cloud. After merging similar and infrequent classes, the dataset includes 16 distinct semantic classes. Implementation Details Prior works experimented mostly with the resolution of $64\times 2048$ for SemanticKITTI [5, 24, 7], and $32\times 960$ [18] or $32\times 2048$ [1] for nuScenes. In contrast, we reduce the azimuth resolution while increasing the projection rate in FLARES mode: resolutions of $64\times 512$ for SemanticKITTI and $32\times 480$ for nuScenes are fixed for the input and the full point cloud is split into up to 3 and 2 ($N_{max}$) sub-clouds during training and inference. We directly use custom configurations of prior works [48, 5, 7, 1] to train the networks in FLARES mode. For training the selected models (excluding RangeViT) on the nuScenes dataset, we standardize the hyperparameter set since no default configurations are provided. Specifically, we use the AdamW optimizer [22] along with a OneCycle scheduler [29], setting the maximum learning rate to $1e^{-3}$ and training for 150 epochs. All models are trained on four NVIDIA GeForce GTX 1080Ti in distributed mode. Evaluation Metrics Following prior works, we assess the performance using Intersection-over-Union (IoU) $\text{IoU}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ and Accuracy (Acc) $\text{Acc}_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$ for each class $i$, and calculate the mean Intersection-over-Union (mIoU) and mean Accuracy (mAcc) across all classes. $TP_{i}$, $FP_{i}$, and $FN_{i}$ represent the true positives, false positives, and false negatives for class $i$, respectively.
4.2 Comparative Study
Groundtruth Baseline FLARES
SalsaNext [7]
Refer to caption Refer to caption Refer to caption
FIDNet [48]
Refer to caption Refer to caption Refer to caption
CENet [5]
Refer to caption Refer to caption Refer to caption
RangeViT [1]
Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative results on SemanticKITTI[2] Points in red and gray represent incorrect and correct predictions, respectively. More examples are provided in the supplementary material.
We compare FLARES with baseline models across two datasets. As shown in Tab. 1, all four networks see significant improvements: SalsaNext has mIoU gains of 5.3% on SemanticKITTI and 2.3% on nuScenes, FIDNet improves by 7.9% and 3.9%, and CENet by 3.3% and 3.1%. RangeViT, as a ViT-based network, also exhibits huge enhancement in performance, confirming FLARES’s generalizability across different architectures. This improvement is especially prominent for smaller, dynamic, and under-represented classes such as truck, motorcycle, bicycle, pedestrian and bicyclist. In Fig. 4, we present a visual comparison. Notably, with the support of FLARES, the network demonstrates improved accuracy in segmenting foreground objects. An exception arises with the motorcyclist class in SemanticKITTI, where IoU scores decrease compared to the baseline. Diving into the problem, this can be traced back to the extremely low occurrence of annotations for that class in the dataset. In standard training on low-resolution range images, this class already suffers from poor representation. In FLARES mode, the occurrence is further reduced by splitting of the point cloud. This accumulation of downsampling prevents the network from optimizing on that rare class effectively and lead to inferior performance. In contrast, the improvement on nuScenes is more consistent as class frequencies are better balanced. We regard this as a corner case when testing on an class-imbalanced dataset. As a future work to resolve the issue, we aim to explore 3D reconstruction techniques to generate real-world-like pseudo LiDAR point clouds for augmentation [4, 23].
In Tab. 2, we further compare FLARES-boosted networks with other state-of-the-arts from various modalities. Given that methods we select preserve relatively fewer number of parameters, FLARES helps to improve the segmentation accuracy, being comparable to other point- or voxel-based approaches that deploy much larger and deeper networks. In addition, they outperform others significantly in latency, achieving superior trade-offs in accuracy and efficiency.
 SemanticKITTI test set
Method
mIoU
car
bicy
moto
truc
o.veh
ped
b.list
m.list
road
park
walk
o.gro
build
fenc
veg
trun
terr
pole
sign
SalsaNext [7] $59.5$ $91.9$ $48.3$ $38.6$ $38.9$ $31.9$ $60.2$ $59.0$ $\mathbf{19.4}$ $91.7$ $63.7$ $75.8$ $\mathbf{29.1}$ $90.2$ $64.2$ $81.8$ $63.6$ $\mathbf{66.5}$ $54.3$ $62.1$
[…]
FIDNet [48] $51.3$ $90.4$ $28.6$ $30.9$ $34.3$ $27.0$ $43.9$ $48.9$ $16.8$ $90.1$ $58.7$ $71.4$ $19.9$ $84.2$ $51.2$ $78.2$ $51.9$ $64.5$ $32.7$ $50.3$
FIDNet $59.5$ $93.9$ $54.7$ $48.9$ $27.6$ $23.9$ $62.3$ $59.8$ $\mathbf{23.7}$ $90.6$ $59.1$ $75.8$ $26.7$ $88.9$ $60.5$ $84.5$ $64.4$ $69.0$ $53.3$ $62.8$
[…]
⋄†CENet [5] $60.7$ $92.1$ $45.4$ $42.9$ $43.9$ $46.8$ $56.4$ $63.8$ $29.7$ $91.3$ $66.0$ $75.3$ $31.1$ $88.9$ $60.4$ $81.9$ $60.5$ $67.6$ $49.5$ $59.1$
CENet $64.7$ $91.9$ $58.6$ $50.3$ $40.6$ $42.3$ $68.9$ $65.9$ $\mathbf{43.5}$ $90.3$ $60.9$ $75.1$ $31.5$ $91.0$ $66.2$ $84.5$ $69.7$ $\mathbf{70.0}$ $61.5$ $67.6$
[…]
RangeViT [1] $64.0$ $95.4$ $55.8$ $43.5$ $29.8$ $42.1$ $63.9$ $58.2$ $\mathbf{38.1}$ $\mathbf{93.1}$ $70.2$ $\mathbf{80.0}$ $32.5$ $\mathbf{92.0}$ $\mathbf{69.0}$ $\mathbf{85.3}$ $\mathbf{70.6}$ $\mathbf{71.2}$ $\mathbf{60.8}$ $64.7$
[…]
 nuScenes val set
Method (year)
mIoU
barrier
bicy
bus
car
const
moto
ped
traffic.c
trailer
truck
driv
o.flat
side
terrain
manm
veg
SalsaNext $72.2$ $74.8$ $34.1$ $85.9$ $88.4$ $42.2$ $72.4$ $72.2$ $\mathbf{63.1}$ $61.3$ $76.5$ $96.0$ $\mathbf{70.8}$ $71.2$ $71.5$ $86.7$ $84.4$
[…]
FIDNet $72.7$ $73.0$ $36.0$ $87.8$ $86.0$ $45.6$ $74.1$ $73.9$ $62.5$ $67.1$ $77.7$ $94.3$ $69.8$ $72.2$ $72.1$ $86.1$ $84.5$
[…]
CENet $73.7$ $73.6$ $32.9$ $92.7$ $87.1$ $\mathbf{53.5}$ $76.1$ $69.0$ $58.7$ $66.8$ $81.6$ $95.6$ $71.1$ $73.7$ $73.2$ $87.5$ $85.7$
[…]
RangeViT $75.2$ $75.5$ $\mathbf{40.7}$ $88.3$ $90.1$ $49.3$ $79.3$ $77.2$ $\mathbf{66.3}$ $65.2$ $80.0$ $96.4$ $71.4$ $73.8$ $73.8$ $\mathbf{89.9}$ $\mathbf{87.2}$
[…]
Table 1: Comparisons of state-of-the-art LiDAR semantic segmentation methods on the test set of SemanticKITTI [2] and val set of nuScenes [3] in standard and  FLARES mode. IoU scores are reported in percentages (%). For each method block, bold and underline indicate the best and second best result in the column. Baseline results trained on low-resolution ($64\times 512$) range images. Models inferred with test-time augmentation [18]. Note that we did not use model ensembling to further boost the model performance.
[…]
PolarNet [46] [’20] 13.6M 71 Polar 71.0 54.9 57.5
SPVNAS [31] [’20] 12.5M 259 Voxel - 64.7 66.4
RandLA-Net [42] [’20] 1.2M 55 Point - - 53.9
Tornado-Net [12] [’20] - - Multiple - 64.5 63.1
[…]
Cylinder3D [49] [’21] 56.3M 170 Voxel 76.1 67.8 65.9
RPVnet [43] [’21] 24.8M 168 Multiple 77.6 68.2 70.3
FPS-Net [40] [’21] 55.7M 48 Range - 54.9 57.1
Lite-HDSeg [28] [’21] - 50 Range - 64.4 63.8
[…]
Meta-RSeg [34] [’22] 6.8M 46 Range - 60.3 61.0
PVKD [15] [’22] 14.1M 76 Voxel 76.0 66.4 71.2
PTv2 [38] [’22] 12.8M 213 Point 80.2 70.3 72.6
2DPASS [45] [’22] 26.5M 119 Multiple 79.4 69.3 72.2
GFNet [27] [’22] - 100 Multiple 76.8 63.2 65.4
[…]
4.3 Ablation Study
To perform the ablation study, we test with CENet [5] on val set of SemanticKITTI [2]. Training Schema As shown in Tab.3, we conduct a comprehensive evaluation to examine the trade-off between accuracy and efficiency under different training configurations. In the standard mode, a high-resolution range image is used as the input to reproduce baseline results. While training in the STR paradigm[18], there is a slight performance drop, the memory consumption is significantly reduced by partitioning the full point cloud through azimuth-wise grouping. The limited increase in latency is due to the fact that STR has to additionally unite the batched prediction along azimuth resolution and the post-processing is still performed on the high-resolution image. In contrast, FLARES achieves remarkable improvements, showing increases of 2.7% in mIoU, 2.1% in mAcc, and 45% acceleration in inference. Similar to STR, FLARES reduces memory consumption compared to the standard mode, further optimizing the efficiency of range-view semantic segmentation.
Method Input resolution mIoU mAcc Lat.
Standard $1\times 64\times 2048$ 64.8 77.2 44 ms
STR [18] $1(5)\times 64\times 480$ 64.3 76.8 41 ms
FLARES $1(3)\times 64\times 512$ 67.5 79.3 24 ms
Table 3: Performance with varying training schemes is evaluated on val set of SemanticKITTI [2]. Input resolution is formatted in $B\times H\times W$ (batch size, image width and height). For STR [18], we use the configuration yielding the best performance on CENet [5]: a full resolution of $64\times 1920$ split into four $64\times 480$ sub-images. For the fair comparison, all models are integrated with proposed components and trained from scratch.
Data Augmentation As presented in Fig. 5(a), various data augmentation methods can enhance the segmentation performance by large margins and our WPD+ performs the best among them, which increases baseline mIoU by 5.7% and mAcc by 2.9%. Mix3D [25] increases the contextual information per frame by fusing one scene to another, however, the model can still suffer from class imbalance. RangeAug [18] consists of 4 different range-wise operations to enrich the semantic and structural cues. From their experiments, it demonstrated that this augmentation technique is especially effective for attention-based networks [33, 9], which possess significantly higher model capacities and are more reliant on data diversity for optimal performance. Similar to their experimental results, applying this method to lightweight CNN-based networks has shown limited success in improvement.
(a) WPD+
(b) #Frames sampled in WPD+
(c) Inclusion of synthetic dataset
(d) MCF
Figure 5: a) Initialization of training with standard geometric data augmentation (GDA) and benchmark several state-of-the-art 3D augmentation techniques, including Mix3D [25], RangeAug [18], and original WPD [13]. b) Different number of sampled frames for best performance c) A comparative plot showing the IoU scores of top-rare classes in scenarios both with and without the inclusion of the synthetic dataset. As reference, the class frequencies in the val set are provided. d) The models are trained using two different input configurations: either a single range image derived from the full cloud (FC) or a range image generated from a sub-cloud (SC), and inferred in FLARES mode. Note that for all trained models in a) - d), we leverage KNN Ensembling during post-processing phase.
WPD+ includes two tunable parameters: the number of sampled frames from the original dataset and the use of the synthetic dataset. To find out the best parameter set, we conduct two additional experiments. Fig. 5(b) shows that sampling 6 frames results in the optimal performance, while increasing the number of frames beyond this point leads to performance degradation. This is likely due to the model’s limited scalability, similar to the effects observed in case of RangeAug[18], in addition, pasting too many pixels belonging to specific classes will again corrupt the semantic balance. Furthermore, Fig. 5(c) illustrates that the synthetic dataset plays a key-role in refining semantic prediction of top-rare classes. Noteworthily, using the synthetic dataset is efficient and practical because this allows us to customize sensor configurations to align with the target dataset and to define specific objects within the scene for downstream applications without any labor cost.
In the next stage, we study how unoccupied pixels in the range image can affect the performance. Firstly, we use the full point cloud for training and sub-clouds for inference, as the decreasing azimuth resolution can already resolve the problem of low 2D occupancy (the first column in Fig. 5(d)). However, empirical results reveal a drop in 2D accuracy when inferring from sub-clouds. This is possibly because of the domain shift caused by the occupancy difference in range images. In the next trial, we use the sub-cloud as training input instead, but this directly limits the performance due to the occupancy reduction. To tackle this compounded challenge brought by FLARES, we introduce Multi-Cloud Fusion, an additional data augmentation step which fuses multiple sub-clouds through occupancy padding during projection phase. As shown in Fig. 5(d), using MCF exhibits the highest performance in the series, achieving 1% increase in IoU over the model trained on full cloud. As an alternative, we apply RangeIP [44], an interpolation-based augmentation technique, to enhance 2D occupancy in the range image, but it results in slight worse accuracy compared to the baseline.
(a) Ablation on Post-Processing
(b) Inference Time [ms]
(c) Ablation on Input Resolution
Figure 6: a) - b): Various post-processing techniques are applied to the same trained model. KNN [24] and NLA [48] are used on the single range image generated from the full point cloud, whereas KNN Ensembling operates on multiple range images derived from sub-clouds. For NNRI, we assess it using both full cloud and sub-clouds. c): Evaluation on different input resolutions and corresponding number of sub-clouds ($N_{max}$).
Post-Processing As introduced in Sec. 3.3, a simple LiDAR model is built to approximate the distance-density function of 3D points and compute cut-off values for valid neighbors extraction. To verify the necessity of the step, some qualitative results are provided in Fig. 7. As can be seen, the adaptive cut-off values refine semantic predictions by better accommodating objects with varying density scales.
(a) Scene
(b) GT
(c) Constant
(d) Adaptive
Figure 7: Segmentation results with different cut-off values in NNRI: in the case of constant value (set at 1), overlapping points of Road are partially misclassified as Car in the top image. Similarly, in the bottom image, half of the points that belong to Building are incorrectly predicted as Vegetation.
Next, we explore the impact of various post-processing techniques on segmentation performance in Fig. 6(a) and 6(b). Regarding conventional KNN [24] as the baseline, NLA [48] demonstrates similar performance in both accuracy and latency. In contrast, we deploy our approach (NNRI) in the standard mode as well and observe a significant improvement: inference time is cut nearly 16% compared to KNN, while mAcc and mIoU increase by 2.8% and 2%, respectively. Unlike KNN, NNRI avoids the computational cost of Gaussian kernel calculations for distance weighting and directly performs nearest neighbor searches on the range image instead of in 3D space, further reducing computational overhead. NNRI interpolates class-wise scores based on relative depths rather than directly voting on hard labels, relying more on weighted information from nearest neighbors, which is the major reason why it outperform other post-processing approaches.
Switching to FLARES mode, we first extend the standard KNN approach to KNN Ensembling, which iteratively gathers votes from all points in each sub-cloud and aggregates them for the final prediction. While this extension improves the accuracy, it comes at approximately doubled latency cost. Conversely, when NNRI is adapted, it consistently provides notable improvements in both efficacy and efficiency. As a reference, we included evaluation scores on 2D predictions (-- dashed lines in Fig. 6(a)), showing that FLARES with NNRI significantly narrows the accuracy gap between 2D and 3D predictions. This suggests that our approach effectively mitigates the ”many-to-one” problem, offering substantial gains in segmentation performance. Input Resolution To further explore the optimal results in FLARES mode, we test various input resolutions, as shown in Fig. 6(c). From the experimental results, we found that resolutions of $512$ and $640$ deliver the best mIoU scores. Increasing the azimuth resolution beyond this point causes a slight performance drop. Nevertheless, all tested configurations outperform the baseline (the first row in Tab.3), demonstrating the superb effectiveness of our approach on range-view LiDAR semantic segmentation.
5 Conclusion
In this work, we introduced FLARES, an optimized training and inference schema designed for seamless integration into any range-view-based network. To enhance the scalability, we further developed two data augmentation techniques for class rebalance and occupancy increase. Additionally, we proposed a novel unsupervised post-processing method to effectively address the ”many-to-one” issue, minimizing the 2D-to-3D performance gap. Our approach has substantiated significant improvements in accuracy and efficiency over baselines across various network architectures on two widely used LiDAR benchmarks. Despite the limitation in some corner case of class imbalance in the dataset, our approach shows its overall versatility and effectiveness in advancing LiDAR semantic segmentation.